OpenStudy (kendricklamar2014):

Write y = -(3/5x) - 2 in standard form using integers. A. 3x + 5y = -10 B. 3x + 5y = -2 C. 3x - 5y = 10 D. -3x + 5y = 10

OpenStudy (texaschic101):

standard form equations do not contain fractions...so we will first multiply the entire equation by the common denominator, which is 5, giving us : 5y = -3x - 10 -- now add 3x to both sides 3x + 5y = -10

OpenStudy (texaschic101):

any time you have fractions and you do not want to mess with them, just multiply entire equation by the common denominator and it will get rid of the fractions

OpenStudy (anonymous):

Standard form is: \[Ax + By = C\] So you would want to start of by moving the X term to the left side by adding it to both sides. \[y + 3/5x = -3/5x + 3/5x -2\] Simplify: [\3/5x + y= -2\] However it needs to be an integer. Multiply by 5: [\3x + 5y = -10\] The answer is A.

OpenStudy (kendricklamar2014):

Ok, I get it, thanks @texaschic101 and @SpooksDF

OpenStudy (texaschic101):

sure thing...thats what we are here for :)

OpenStudy (anonymous):

No problem.